# Coupled dimerized alternating-bond quantum spin chains in the distorted honeycomb-lattice magnet Cu5SbO6

### v1

@@ III.1 Crystal structure and stacking faults
A selected-area electron diffraction (SAED) pattern for the powder sample synthesized at 1000 C is shown in Fig. 3(d) in which diffusive scattering streaks are observed along $(2,0,L)$ and $(4,0,L)$ but are absent along $(0,0,L)$ and $(6,0,L)$. The streaks observed reflect lack of periodicity in the $c^{*}$ direction pointing to translations of the Cu2+ honeycomb layers that intermittently intervene the ideal ‘123’ sequence. To gain more insight into this mis-translation, we next investigate a TEM image of the $ac$ plane of this sample as displayed in Fig. 3(c). The image clearly features an irregularity in the stacking pattern and hints at the presence of two in-plane translations along the $a$ direction. Inspecting Fig. 3(b) suggests that, in addition to the ideal in-plane translation, there is another possible translation (depicted by purple arrows) of the Cu2+ honeycomb layers that translates Sb5+ from sub-lattice ‘1’ to ‘3’, ‘3’ to ‘2’, and ‘2’ to ‘1’. Thus, an appropriate model for stacking faults in Cu5SbO6 corresponds to intervening the ideal ‘123’ sequence by a stacking-fault translation at certain frequency. It is worth noting that the ideal in-plane translation vector can be represented by (1/3, 0, 0) and that of the stacking fault by $(2/3,0,0)$ or, equivalently, (1/3,0,0)(-1/3,0,0). This model immediately provides a description for the disappearance of diffusive streaks along $(6,0,L)$. As the stacking-fault translation vector is $(2/3,0,0)$ which can be obtained by adding $(1/3,0,0)$ to the ideal translation vector. The additional phase factor caused by the extra $(1/3,0,0)$ results in a phase difference of $2\pi$ for $(6,0,L)$ planes, hence, constructive interference. The streaks are not observed along $(0,0,L)$ since the phase factor due to these planes always adds in-phase irrespective of the in-plane translations. To further verify the model, a SAED pattern is simulated using faults software package [Treacy1991, Casas-Cabanas2006, Casas-Cabanas2016]. The frequency of the ideal $p_{1}$ and of the stacking-fault translations $p_{2}$ are introduced which are subjected to a constraint $p_{1}+p_{2}=1$. Fig. 3(f) illustrates a simulated SAED pattern of an extreme case where $p_{1}=p_{2}=1/2$ which can well reproduce the diffusive intensities along $(2,0,L)$ and $(4,0,L)$ in the observed SAED pattern in Fig. 3(d). In addition, the model can also reproduce the broad peak in the XRD data of a powder sample as illustrated in the inset of Fig. 5(a) of Ref. [Climent2011].
@@
A SAED pattern for the sample synthesized at 1150 C, a temperature just below the melting point, shows a strikingly different result as shown in Fig. 3(e). In this case, the diffusive scattering streaks are unobserved reflecting a restoration of the periodicity along the $c^{*}$ direction. The absence of the diffusive streaks in this SAED pattern is consistent with the absence of peak broadening, especially in $19^{\circ}<2\theta<23^{\circ}$, in the powder XRD data collected from ground crystals (see Fig. 2(b)). The improvement of stacking faults in Cu5SbO6 by raising the synthesis temperature was also reported in Ref. [Climent2011].
@@ III.2 Magnetization
We perform quantum Monte Carlo (QMC) simulations to acquire a model for $\chi_{\mathrm{dim}}(T)$. The reduced susceptibility, $\chi^{*}(t)$, for $0.01\leq t\leq 10$ where $t=k_{\mathrm{B}}T/J_{\mathrm{max}}$ is calculated using the loop algorithm [Todo2001, Evertz2003] implemented in the alps simulation package [Bauer2011]. The simulation is performed for the $J_{1}$-$J_{2}$-$J_{4}$ model with $J_{\mathrm{max}}=J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV. The determination of these coupling parameters will be discussed in the next section. The simulations are performed on a cluster of $\times100$ spins with a periodic boundary condition using 100 000 Monte Carlo steps for thermalization and 500 000 Monte Carlo steps after thermalization. To obtain a functional form of $\chi^{*}(t)$, we fit simulation results with [Johnston2000]
@@
To reaffirm the optimized parameters, we further perform quantum Monte Carlo (QMC) simulations to calculate the reduced magnetization, $M^{*}(h)$, with the same set of coupling parameters for $0\leq h\leq 7\times 10^{-2}$ where $h=g\mu_{\mathrm{B}}H/J_{\mathrm{max}}$ and $T=120$ K using the directed-loop [Syljuasen2002, Syljuasen2003] algorithm. The calculation conditions are the same as for $\chi^{*}(t)$. $M_{\mathrm{dim}}(H)$ is related to $M^{*}(h)$ by
@@
The simulated $M_{\mathrm{dim}}(H)$ is calculated with the fitted Landé factor $g$ and then compared with the background-corrected experimental data. As seen in Figs. 4(b) and 4(c), a reasonable agreement is achieved between the data and model calculations confirming the validity of the Landé factors and the exchange couplings.
@@ III.3 Triplet excitations
Figure 5: Powder averaged TOF INS intensity maps as a function of energy transfer and wavevector at (a) $T=4$ K and (b) $T=200$ K. Yellow dashed lines in (a) mark $\hbar\omega=$ 11 meV and 21 meV indicating the band of magnetic scattering. Solid symbols in (a) are excitation energies extracted from BT-7 spectrometer taken at 2.8 K. Error bars in (a) represent five standard deviations. Density of states (DOS) of (c) the $J_{1}$-$J_{2}$-$J_{3}$ model, and of (d) the $J_{1}$-$J_{2}$-$J_{4}$ model.
@@
Data reduction and powder averaging of the TOF INS data collected from the AMATERAS spectrometer are done with use of the utsusemi software package [Inamura2013]. The obtained intensity maps at 4 K and 200 K are displayed in Figs. 5(a) and 5(b), respectively. Also included in Fig. 5(a) is the excitation energy (circle symbols) extracted from $E$-scan data taken from the BT-7 spectrometer at 2.8 K. To appreciate the overall excitation characteristics of Cu5SbO6, let us compare the intensity maps at the two temperatures. At $T=200$ K where the magnetic correlations completely subside (see Fig. 6(a)), there remain excitations whose intensity noticeably increases with the magnitude of wavevector $Q$ and with temperature. Based on this observation, one can immediately attribute the excitations at 200 K to phonons. Therefore, the excitations observed at $T=4$ K in $1\;\mathrm{\AA }^{-1}\lesssim Q\lesssim 3\;\mathrm{\AA }^{-1}$ are magnetic in nature, as the intensity decreases with increasing $Q$ and $T$. The magnetic excitation band features two peaks in intensity (henceforth $M_{1}$ and $M_{2}$, respectively) at 15 meV and 18 meV and covers an extensive range from the band bottom at $\sim$ 11 meV to the band top at $\sim$ 21 meV amounting to a sizable bandwidth of $\sim$ 10 meV. No more magnetic excitation is detected for $\hbar\omega\lesssim 11$ meV.
@@
To explore the temperature evolution of the excitation spectrum, we next examine the $E$-scan data collected at base and elevated temperatures from the BT-7 spectrometer. Fig. 6(a) shows the data measured at $2.8\;\mathrm{K}\leq T\leq 200\;\mathrm{K}$ with a fixed wavevector $Q=1.96\;\mathrm{\AA }^{-1}$. At $T=2.8$ K, the data exhibit three excitation peaks at 12 meV, 15 meV, and 18 meV, denoted by $P$, $M_{1}$, and $M_{2}$, respectively. At elevated temperatures, $P$ peak becomes stronger whereas $M_{1}$ and $M_{2}$ peaks weaken and disappear at 200 K consistent with the TOF data. We fit the spectra to a Lorentzian function combined with a linear background to acquire the excitation energy, $\hbar\omega$, for each $T$ and $Q$. The width of $M_{1}$ and $M_{2}$ peaks are assumed identical. Fitting the data (i) at $T=2.8$ K and with $1.12\;\mathrm{\AA }^{-1}\leq Q\leq 3.09\;\mathrm{\AA }^{-1}$ and (ii) at $2.8\;\mathrm{K}\leq T\leq 200\;\mathrm{K}$ with $Q=1.96\;\mathrm{\AA }^{-1}$ yields $\hbar\omega$ plotted in Fig. 5(a) as circle symbols and $T$-dependent $\hbar\omega$ shown in Fig. 6(b). In alignment with $\chi(T)$ data, the monotonic increase in $\hbar\omega(T)$ points out that the gapped excitation observed in Cu5SbO6 is a triplet excitation (or triplon) since, in a dimerized magnet, triplons are subjected to the hard-core constraint that prevents them from occupying the same dimer site resulting in repulsion between them [Xu2007, Essler2008, Exius2010, Normand2011] as opposed to long-range ordered magnets for which the gap decreases with increasing temperature due to the reduction in ordered moment.
To acquire an appropriate spin Hamiltonian for Cu5SbO6, we employ the first-order dimer expansion to derive the energy dispersion, $\Delta E_{\boldsymbol{q}}$, of triplet excitations for both $J_{1}$-$J_{2}$-$J_{3}$ and $J_{1}$-$J_{2}$-$J_{4}$ models (a detailed derivation can be found in the appendix). Let us begin with the former model whose dispersion is expressed in Eq. (11). Since a direct comparison of this result with the averaged INS data in Fig. 5(a) is impracticable, we resort to the density of states (DOS), $D(\omega)$, as defined by
@@
where the integration is over a unit cell of the reciprocal lattice formed by dimers. Approximating the delta function by a Lorentzian function, $\delta(x)=\epsilon/\pi(x^{2}+\epsilon^{2})$, where $\epsilon$ is a small positive number, we finally obtain DOS for the $J_{1}$-$J_{2}$-$J_{3}$ model as shown in Fig. 5(c). The result is presented for four different values of the inter-chain coupling $J_{3}$ with $J_{1}=16.95$ meV, $J_{2}=-3.9$ meV. When $J_{3}=0$, two peaks in DOS appear at the top and bottom of the band. As $J_{3}$ increases, the peak at the top shifts towards lower energy and the tail develops towards higher energy. Although $J_{3}\approx 1$ meV correctly reproduces peaks in DOS at 15 meV and 18 meV, this model does not provide a tail below 15 meV regardless of the value of $J_{3}$ which is in contradiction to the observed spectrum. Therefore, the $J_{1}$-$J_{2}$-$J_{3}$ model can be ruled out.
We next turn to the $J_{1}$-$J_{2}$-$J_{4}$ model. To obtain DOS for this model, we first consider the dispersion in Eq. (12) from which the energy of $M_{1}$ and $M_{2}$ peaks, denoted by $E_{1}$ and $E_{2}$, respectively, are given by
@@
which give $2J_{1}=E_{1}+E_{2}$ and $|J_{2}|=|J_{4}|+(E_{2}-E_{1})$. Using $E_{1}=15.0(1)$ meV and $E_{2}=18.0(2)$ meV, we can identify $J_{1}=16.5(1)$ meV and $|J_{2}|=|J_{4}|+3.0$ meV. Fig. 5(d) shows DOS of the $J_{1}$-$J_{2}$-$J_{4}$ model calculated with four different values of $|J_{4}|$. With these coupling parameters the DOS correctly reproduces peaks at 15 meV and 18 meV. Moreover, in contrast to the previous model, as $J_{4}$ increases, tails in DOS develop at both higher and lower energy which resembles the pattern observed in the INS spectrum. Based on the INS data, one can obtain an estimate of $|J_{4}|\approx 3$ meV, and, thus, $|J_{2}|\approx 6$ meV. It is worth noting that in the presence of stacking faults, although the magnitude of $J_{4}$ might be unchanged as the path responsible for the coupling still involves the same ions and geometry, the dimer connectivity is different and the $J_{1}$-$J_{2}$-$J_{4}$ model might be inapplicable. However, based on the XRD result shown in Fig. 2(b), the amount of stacking faults in our sample employed to perform INS experiment, if any, is minimal.

### v2

@@ III.1 Crystal structure and stacking faults
A selected-area electron diffraction (SAED) pattern for the powder sample synthesized at 1000 C is shown in Fig. 3(d) in which diffusive scattering streaks are observed along $(2,0,L)$ and $(4,0,L)$ but are absent along $(0,0,L)$ and $(6,0,L)$. The streaks observed reflect lack of periodicity in the $c^{*}$ direction pointing to translations of the Cu2+ honeycomb layers that intermittently intervene the ideal ‘123’ sequence. To gain more insight into this mis-translation, we next investigate a TEM image of the $ac$ plane of this sample as displayed in Fig. 3(c). The image clearly features an irregularity in the stacking pattern and hints at the presence of two in-plane translations along the $a$ direction. Inspecting Fig. 3(b) suggests that, in addition to the ideal in-plane translation, there is another possible translation (depicted by purple arrows) of the Cu2+ honeycomb layers that translates Sb5+ from sub-lattice ‘1’ to ‘3’, ‘3’ to ‘2’, and ‘2’ to ‘1’. Thus, an appropriate model for stacking faults in Cu5SbO6 corresponds to intervening the ideal ‘123’ sequence by a stacking-fault translation at certain frequency. It is worth noting that the ideal in-plane translation vector can be represented by (1/3, 0, 0) and that of the stacking fault by $(2/3,0,0)$ or, equivalently, (1/3,0,0)(-1/3,0,0). This model immediately provides a description for the disappearance of diffusive streaks along $(6,0,L)$. As the stacking-fault translation vector is $(2/3,0,0)$ which can be obtained by adding $(1/3,0,0)$ to the ideal translation vector. The additional phase factor caused by the extra $(1/3,0,0)$ results in a phase difference of $2\pi$ for $(6,0,L)$ planes, hence, constructive interference. The streaks are not observed along $(0,0,L)$ since the phase factor due to these planes always adds in-phase irrespective of the in-plane translations. To further verify the model, a SAED pattern is simulated using faults software package [60, 61, 62]. The frequency of the ideal $p_{1}$ and of the stacking-fault translations $p_{2}$ are introduced which are subjected to a constraint $p_{1}+p_{2}=1$. Fig. 3(f) illustrates a simulated SAED pattern of an extreme case where $p_{1}=p_{2}=1/2$ which can well reproduce the diffusive intensities along $(2,0,L)$ and $(4,0,L)$ in the observed SAED pattern in Fig. 3(d). In addition, the model can also reproduce the broad peak in the XRD data of a powder sample as illustrated in the inset of Fig. 5(a) of Ref. [48].
@@
A SAED pattern for the sample synthesized at 1150 C, a temperature just below the melting point, shows a strikingly different result as shown in Fig. 3(e). In this case, the diffusive scattering streaks are unobserved reflecting a restoration of the periodicity along the $c^{*}$ direction. The absence of the diffusive streaks in this SAED pattern is consistent with the absence of peak broadening, especially in $19^{\circ}<2\theta<23^{\circ}$, in the powder XRD data collected from ground crystals (see Fig. 2(b)). The improvement of stacking faults in Cu5SbO6 by raising the synthesis temperature was also reported in Ref. [48].
@@ III.2 Magnetization
We perform quantum Monte Carlo (QMC) simulations to acquire a model for $\chi_{\mathrm{dim}}(T)$. The reduced susceptibility, $\chi^{*}(t)$, for $0.01\leq t\leq 10$ where $t=k_{\mathrm{B}}T/J_{\mathrm{max}}$ is calculated using the loop algorithm [63, 64] implemented in the alps simulation package [65]. The simulation is performed for the $J_{1}$-$J_{2}$-$J_{4}$ model with $J_{\mathrm{max}}=J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV. The determination of these coupling parameters will be discussed in the next section. The simulations are performed on a cluster of $\times100$ spins with a periodic boundary condition using 100 000 Monte Carlo steps for thermalization and 500 000 Monte Carlo steps after thermalization. To obtain a functional form of $\chi^{*}(t)$, we fit simulation results with [66]
@@
To reaffirm the optimized parameters, we further perform quantum Monte Carlo (QMC) simulations to calculate the reduced magnetization, $M^{*}(h)$, with the same set of coupling parameters for $0\leq h\leq 7\times 10^{-2}$ where $h=g\mu_{\mathrm{B}}H/J_{\mathrm{max}}$ and $T=120$ K using the directed-loop [67, 68] algorithm. The calculation conditions are the same as for $\chi^{*}(t)$. $M_{\mathrm{dim}}(H)$ is related to $M^{*}(h)$ by
@@
The simulated $M_{\mathrm{dim}}(H)$ is calculated with the optimized Landé factor $g$ and then compared with the background-corrected experimental data. As seen in Figs. 4(b) and 4(c), a reasonable agreement is achieved between the data and model calculations confirming the validity of the Landé factors and the exchange couplings.
@@ III.3 Triplet excitations
Figure 5: (a) The observed powder-averaged magnetic intensity map at $T=4$ K (b) The calculated powder-averaged magnetic intensity map using the $J_{1}$-$J_{2}$-$J_{4}$ model with $J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV. (c) The density of states (DOS) of the $J_{1}$-$J_{2}$-$J_{3}$ model and of (d) the $J_{1}$-$J_{2}$-$J_{4}$ model.
@@
Data reduction and powder averaging of the TOF INS data collected from the AMATERAS spectrometer were performed using the utsusemi software package [52]. The obtained intensity maps at $T=4$ K and 200 K are displayed in Figs. S.1(a) and S.1(b), respectively, in the Supplemental Information (SI). Also included in Fig. S.1(a) is the excitation energy (circle symbols) extracted from the $E$-scan data taken from the BT-7 spectrometer at $T=2.8$ K. In Fig. S.1(b), at $T=200$ K where the magnetic correlations are expected to subside (see also Fig. 6(a)), there remain excitations whose intensity noticeably increases with the magnitude of the wavevector $Q$ and temperature. We, therefore, attribute these excitations to phonons. To obtain the magnetic scattering intensity map, $I_{\mathrm{mag}}(Q,\omega)$, as shown in Fig. 5(a), we subtracted the data at the two temperatures after rescaling the data at 200 K by the temperature-dependent factor $1+n(T)=(1-\exp(-E/k_{\mathrm{B}}T)^{-1}$ where $n(T)$ is the Bose factor. It should be noted that the scattering intensity seen at $\hbar\omega=12$ meV and $Q>3$ Å-1 is most likely due to residual phonon intensity. To acquire the key characteristics of the magnetic scattering, $I_{\mathrm{mag}}(Q,\omega)$ is then integrated over the wavevector $1\;\mathrm{\AA ^{-1}}\leq Q\leq 3\;\mathrm{\AA ^{-1}}$. The integrated intensity shown in Fig. S.2(b) features two peaks at $\hbar\omega=15$ meV and 18 meV and covers an extensive range from the band bottom at $\sim$ 11 meV to the band top at $\sim$ 21 meV amounting to a sizable bandwidth of $\sim$ 10 meV. No magnetic excitation is detected at $\hbar\omega\lesssim 11$ meV or $\hbar\omega\gtrsim 21$ meV.
@@
To explore the temperature evolution of the excitation spectrum, we next examine the TA $E$-scan data collected at base and elevated temperatures from the BT-7 spectrometer. Fig. 6(a) shows the data measured at $2.8\;\mathrm{K}\leq T\leq 200\;\mathrm{K}$ with a fixed wavevector $Q=1.96\;\mathrm{\AA }^{-1}$. At $T=2.8$ K, the data exhibit three excitation peaks at 12 meV, 15 meV, and 18 meV, denoted by $P$, $M_{1}$, and $M_{2}$, respectively. At elevated temperatures, the $P$ peak becomes stronger whereas the $M_{1}$ and $M_{2}$ peaks weaken and eventually disappear at 200 K. Therefore, the $M_{1}$ and $M_{2}$ peaks are consistent with the two peaks observed in the integrated magnetic intensity shown in Fig. S.2(b). For the $P$ peak, however, it is essential to note that in the TOF data (see Figs. S.1(a) and S.1(b) in the SI) there exist a phonon band as well as magnetic scattering at $\hbar\omega=12$ meV. Therefore, the $P$ peak, especially at $T=2.8$ K, results from the superposition of these two sources. As the temperature is elevated, the magnetic contribution diminishes leaving only the enhanced phonon contribution. We fit the spectra to a Lorentzian function combined with a linear background to acquire the excitation energy, $\hbar\omega$, for each $T$ and $Q$. The width of the $M_{1}$ and $M_{2}$ peaks are assumed identical. Fitting the data (i) at $T=2.8$ K and with $1.12\;\mathrm{\AA }^{-1}\leq Q\leq 3.09\;\mathrm{\AA }^{-1}$ and (ii) at $2.8\;\mathrm{K}\leq T\leq 200\;\mathrm{K}$ with $Q=1.96\;\mathrm{\AA }^{-1}$ yields $\hbar\omega$ plotted as circle symbols in Fig. S.1(a) and a $T$-dependent $\hbar\omega$ shown in Fig. 6(b). In alignment with $\chi(T)$ data, the monotonic increase in $\hbar\omega(T)$ points out that the gapped excitation observed in Cu5SbO6 is a triplet excitation (or triplon) since, in a dimerized magnet, triplons are subjected to the hard-core constraint that prevents them from occupying the same dimer site resulting in repulsion between them [69, 70, 71, 72] as opposed to long-range ordered magnets for which the gap decreases with increasing temperature due to the reduction in ordered moment.
To acquire an appropriate spin Hamiltonian for Cu5SbO6, we employ the first-order dimer expansion to derive the energy dispersion, $\Delta E_{\boldsymbol{Q}}$, of triplet excitations for both $J_{1}$-$J_{2}$-$J_{3}$ and $J_{1}$-$J_{2}$-$J_{4}$ models (a detailed derivation can be found in the Appendix). The expressions for the dispersion of both models are given in Eqs. (14) and (15). In these expressions, the wavevector $\boldsymbol{Q}$ is written in terms of the basis vectors $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$, and $\boldsymbol{b}_{3}$ defined in Eqs. (8)-(10). Here, $q_{i}$ represents the component of each basis vector. Let us begin with the $J_{1}$-$J_{2}$-$J_{3}$ model with the dispersion given in Eq. (14). For an isolated FM-AFM chain, the dispersion along the $q_{1}$ direction oscillates around the energy set by the intra-dimer AFM $J_{1}$ coupling with the amplitude determined by the magnitude of the inter-dimer FM $J_{2}$ coupling. The incorporation of the interchain AFM $J_{3}$ coupling predominantly only results in a modulation of the top of the dispersion along the $q_{2}$ direction (see Figs. 8(a)-8(d)). To see how the dispersion produces the $M_{1}$ and $M_{2}$ peaks in the magnetic scattering, we consider the density of states (DOS), $D(\omega)$, as defined by
@@
where the integration is over a unit cell of the reciprocal lattice formed by dimers. Approximating the delta function by a Lorentzian function, $\delta(x)=\epsilon/\pi(x^{2}+\epsilon^{2})$, where $\epsilon$ is a small positive number, we finally obtain DOS for the $J_{1}$-$J_{2}$-$J_{3}$ model as shown in Fig. 5(c). The result is presented for four different values of the interchain coupling $J_{3}$ with $J_{1}=16.95$ meV, $J_{2}=-3.9$ meV. When $J_{3}=0$, two peaks appear at the top and bottom of the DOS. As $J_{3}$ increases, the peak at the top shifts towards lower energy and the tail develops towards higher energy. Although $J_{3}\approx 1$ meV correctly reproduces peaks in DOS at 15 meV and 18 meV, this model does not provide a tail below 15 meV regardless of the value of $J_{3}$ which is in contradiction to the observed spectrum. Therefore, the $J_{1}$-$J_{2}$-$J_{3}$ model is not viable in explaining the data.
We next turn to the $J_{1}$-$J_{2}$-$J_{4}$ model. The introduction the interlayer AFM $J_{4}$ coupling to the system of isolated FM-AFM chains delivers a slightly distinct effect on the dispersion. Unlike with $J_{3}$, it is found that modulation along the $q_{3}$ direction appears at both the top and the bottom of the dispersion (see Figs. 8(e)-8(h)). To obtain DOS for this model, we first consider the dispersion in Eq. (15) from which the energy of $M_{1}$ and $M_{2}$ peaks, denoted by $E_{1}$ and $E_{2}$, respectively, are given by
@@
which give $2J_{1}=E_{1}+E_{2}$ and $|J_{2}|=|J_{4}|+(E_{2}-E_{1})$. Using $E_{1}=15.0(1)$ meV and $E_{2}=18.0(2)$ meV, we can identify $J_{1}=16.5(1)$ meV and $|J_{2}|=|J_{4}|+3.0$ meV. Fig. 5(d) shows DOS of the $J_{1}$-$J_{2}$-$J_{4}$ model calculated with four different values of $|J_{4}|$. With these coupling parameters the DOS correctly reproduces peaks at 15 meV and 18 meV. Moreover, in contrast to the previous model, as $J_{4}$ increases, tails in DOS develop at both higher and lower energy which resembles the pattern observed in the INS spectrum. Based on the INS data, one can obtain an estimate of $|J_{4}|\approx 3$ meV, and, thus, $|J_{2}|\approx 6$ meV. It is essential to note that in the presence of stacking faults, although the magnitude of $J_{4}$ might be unchanged as the path responsible for the coupling still involves the same ions and geometry, the dimer connectivity is different and the $J_{1}$-$J_{2}$-$J_{4}$ model might be inapplicable.
It is natural to extend the model to include the $J_{3}$ coupling in the $J_{1}$-$J_{2}$-$J_{4}$ model. To this end, let us examine the effect of $J_{3}$ on the current model by calculating the DOS of the extended model with different values of $J_{3}$ as described in the SI. As can be seen in Fig. S.2(c), the presence of the $J_{3}$ coupling results in a suppression of the 18 meV peak and a further elongation of the tail above 21 meV. These effects become more enhanced as $J_{3}$ grows. Comparison with the integrated magnetic intensity in Fig. S.2(b) suggests an upper bound for $J_{3}$ of merely $\sim 0.3$ meV, which can safely be neglected when taking the uncertainty of the coupling parameters into account. Finally, to verify the $J_{1}$-$J_{2}$-$J_{4}$ model, a powder-averaged INS spectrum was simulated using the proposed set of coupling parameters as outlined in the SI. The simulated result shown in Fig. 5(b) compares very well with the observed spectrum justifying the $J_{1}$-$J_{2}$-$J_{4}$ model as adequate and appropriate for Cu5SbO6.
Future INS studies on high-quality single-crystal samples may shed more light on some exotic aspects of the triplon excitations and allow the spin-spin correlations to be mapped out more precisely. The latter, as discussed in Refs. [73, 74, 75], play a crucial role in determining the degree of hidden quantum entanglement in a many-body system such as a spin chain. It has been argued in these studies that the spatial extent of quantum entanglement is directly related to the spin-spin correlations that can be measured experimentally by means of the neutron scattering technique. In this arena, as pointed out in Ref. [22], an alternating FM-AFM Heisenberg spin chain may triumph over its alternating AFM-AFM counterpart owing to the non-zero spin-spin correlations on both FM and AFM bonds. Given this, Cu5SbO6 can serve as a model system for exploring the spatial evolution of quantum entanglement manifesting in a many-body system.